# Discovering Properties of Inflectional Morphology in Neural Emergent Communication

### v1

@@ 5.2 Natural Languages
As a comparison to humanlike communication schemes, we also analyze subsets of the verb conjugations of two natural languages, Spanish and Arabic, as Attr-Val languages. We chose these due to the availability of high-quality datasets, the fact that both conjugate for tense and person, and the fact that they represent fairly divergent communication schemes along the concatenativity axis. Spanish data comes from the Jehle Spanish Verb Database Jehle (2011), and Arabic data was sourced from Nawaz et al. (2025). We select a set of verb attributes comparable to the inflection Attr-Val setting: root, tense, and person. We then generate Attr-Val configurations based on combinations of these verb attributes sampled from the data, which serve as cognitively plausible (albeit simplified) representations of actual cognitive objects and events encoded and reconstructed in natural communication. For other attributes beyond tense and person, we choose a constant value (e.g., mood is always indicative). We generate 50 such inflectional sublanguages for Spanish and Arabic, each of which are mappings between lexeme-slot tuples and surface forms , as per Wu et al. (2019).
We find that up to 76% of generated Spanish sublanguages can be meaningfully segmented by either HAS or BPE (i.e., BoSDis ratio greater than 1), while only up to 8% of the Arabic sublanguages can be.888Detailed analysis in Appendix E. The latter demonstrates how both tools for symbol segmentation are insufficient for nonconcatenative morphology.
6 EmCom Experimental Results
Our reimagined Attr-Val settings targeting double articulation and inflection, general and specific evaluation metrics for communication schemes, and illustrative artificial and natural language reference points enable us to ask previously unexplored research questions around morphology in EmCom. To begin this exploration, we specifically ask:
@@ 6 EmCom Experimental Results
Do emergent languages exhibit fusionality in inflection similarly to natural languages?
In our experiments, agents are trained with eight different random initializations in the default and inflection Attr-Val games. Nearly all listeners exceed 90% accuracy on the task; more details are provided in Appendix F. Five of seven successful languages exhibit a BoSDis ratio greater than 1 under BPE96 in the default Attr-Val setting, while all do in the inflection setting. These results imply the general existence of multi-character meaningful segments in the emerged languages, confirming that double articulation is achieved.
6.1 Concatenativity and Ease of Articulation
@@
Prior work has often explored pressures inspired by human evolution, such as ease of learning Kirby et al. (2014); Li and Bowling (2019) among others Vithanage et al. (2023); Galke and Raviv (2024), to encourage compositionality in emergent communication schemes. Meanwhile, prior work often assumes or favors concatenativity in their evaluations of schemes Resnick et al. (2020); Ueda et al. (2023) despite the possibility of compositional but nonconcatenative communication schemes (as demonstrated in our artificial and natural languages).
@@
In order to achieve a goal of concatenative emergent languages with double articulation, we hypothesize that an appropriate pressure to explore may be ease of articulation. Human spoken languages are universally subject to phonological rules which drive the possibility of vocal sounds based on the environment and vocal tract physiology Dediu et al. (2017). The interface between phonology and morphology is a key area of study in linguistics, and a computational evolutionary model of morphology (as we hope to lay the groundwork for) would be incomplete without considering that morphemes are made up of meaningless components subject to similar rules independent of their meanings. Specifically, there are disallowed clusters of sounds in spoken languages (e.g., Hawaiian disallows consonant clusters), which may represent a significant pressure towards concatenative languages. To test this, we added a term to the training loss for a toy simulated phonology for agents, where even-valued characters can not occur next to even-valued characters, and odd-valued characters can not occur next to odd-valued characters.999More details provided in Appendix H.
@@
Both without and with this pressure, we visualize the concatenativity (measured by BPELen) of emergent communication schemes in the default and inflection Attr-Val settings respectively in Figures 3 and 4. First, we observe no inductive bias toward concatenavity in our RNN-based agents as previous works implicitly assume; instead, emergent languages are distributed along a spectrum between perfectly concatenative and nonconcatenative. This may not be unrealistic: the natural languages also exhibit such a spectrum. Spanish appears relatively diverse, likely due to the presence of vowel patterns and auxilliary verbs. Arabic is more bimodal, with some very nonconcatenative sublanguages and some which have a high maximal compression due to being very fusional.
@@
However, the introduction of articulation pressure sharply decreases BPELen in emergent languages. We perform additional -tests on the mean BPELen96 of emergent languages between conditions, observing statistical significance in this change for both the default and inflection Attr-Val setting ( and respectively). This suggests ease of articulation is indeed a strong pressure towards concatenation, and the presence of “unpronounceable” clusters at the phonological (character) level is sufficient to induce concatenation at the morphological (symbol) level. More broadly, this may help explain the dominance of concatenative morphology in natural languages.101010In Appendix J, we provide a thorough analysis of concatenativity and fusionality metrics with and without this pressure in both the default and inflection Attr-Val settings.
6.2 TopSim in Inflectional Languages
@@
To shed more light on the compositionality of naturalistic communication schemes, we compare the TopSim of emergent and natural languages in the inflection Attr-Val setting in Figure 5, including emergent languages with the articulation pressure. Compared to natural languages, TopSim for emergent languages was low, except for one instance with a TopSim of .507 (in the range of natural languages). Interestingly, the articulation pressure resulted in a small decrease in TopSim, including one outlier language which had negative TopSim despite exceeding 90% reconstruction accuracy. Together, this suggests that emergent languages still have room for improvement in achieving humanlike topographic similarity, and there may remain more pressures toward this to identify and explore.
6.3 Fusionality Under Inflection
@@
When reinterpreting Attr-Val as inflection, fusionality becomes a naturalistic trait for languages. Suppletion (fusion of a root with grammatical features, e.g., in went as the past tense form of go) is much rarer than fusion of two or more grammatical features. To judge whether this holds in EmCom, we identify the most fused attribute pair which exhibits the greatest F-TopSim in emergent and natural languages. As expected, attributes 1 and 2 (tense and person, which have the lowest number of values) had the highest F-TopSim in 6 of 8 emergent languages, suggesting these attributes are most likely to fuse. This mirrors the fusion of fixed grammatical features in natural language. In Arabic, most sublanguages (78%) also favored the fusion of these attributes. Unexpectedly, in Spanish, attributes 0 (root) and 1 (tense) fused most often (48% of sublanguages), with (1,2) at 30%. This may be due to the fact that specific roots can influence the form of inflectional affixes, e.g., the -amos suffix only applies to -ar verbs like llorar. An explanation for these trends may be that symbolic complexity Zhang (2025) increases the least in fusing these attributes, i.e., only 5 symbols are required to communicate the values for these two attributes before fusion, 6 are required after fusing them. Meanwhile, fusing the root attribute (with 42 values) with another attribute can drastically increase the symbolic complexity (i.e., 2-3 times).
@@ 7 Conclusion
In this work, we outline an experimental paradigm for emergent morphology, motivated by the principle of double articulation and a rich analogy between the Attr-Val game and inflectional morphology. We review existing metrics for EmCom, proposing a new metric for concatenativity based on BPE, and a new metric for fusionality based on variations in TopSims under fusion of attributes.
In our empirical analyses, we calibrate our metrics towards a wide array of idealized compositional schemes that are possible under this paradigm, as well as schemes based on natural language inflection. We then demonstrate through EmCom experiments the power of this reimagined setting to produce linguistically interesting analyses grounded in comparison to natural language. In particular, we show that the double articulation aspect of our paradigm reveals simulated articulatory constraints as a promising pressure toward concatenativity, which prior work has often implicitly assumed in evaluations of compositionality and attempted to elicit through other pressures, and future work may use to test more specific morphophonological hypotheses. Further, our emergent languages rederive the bias towards fusion of grammatical features in natural language. We hope that future work will adopt the small-vocabulary and inflectional Attr-Val settings for focused, linguistically motivated studies of morphology, and will benefit from being able to directly compare to natural language.
@@
In our natural language reference points, we only considered Arabic and Spanish, which are not a representative sample of the world’s natural languages, and were chosen in part due to the availability of data. Without a much larger and more diverse dataset, we cannot definitely say whether a given emergent language is naturalistic. Additionally, the Attr-Val scheme we focused on is a rather small subset of the complexity of natural morphologies. A full analysis of a large set of natural languages, to cover this diversity of morphologies, is outside the scope of this paper, but we believe our examples will provide a solid initial point of comparison for emergent languages. Boldt and Mortensen (2024a) explored larger-scale cross comparisons of emergent languages, and we anticipate that the addition of parallel natural language data to such analyses would prove useful for future work.
@@
8 Acknowledgements
We thank our anonymous reviewers for their thoughtful and constructive feedback on this work in progress. We thank Zheyuan Zhang for helpful discussion about this work, and we thank Ryo Ueda for consulting with us on the implementation of metrics in Ueda et al. (2023).
Appendix A Harris’ Articulation Scheme
We replicated the results in Ueda et al. (2023) that C-TopSim is greater than W-Topsim. This result led the authors to suggest that Criterion 3 for Harris’ Articulation Scheme. (i.e. W-TopSim should be higher than C-TopSim) may be unsolvable.
@@ Appendix A Harris’ Articulation Scheme
Appendix B Positional Disentanglement
@@
Appendix C Artificial Languages Extended Analysis
@@
First, the level of concatenation in the first three artificial languages has insignificant effect on TopSim and learnability, implying a lack of inductive bias toward concatenation in recurrent listener neural networks. Notably, this is not consistent with human results in linguistics Finley and Newport (2021), and implies additional pressures may be required to induce concatenative morphology. Meanwhile, as expected, BoSDis metrics indicate that both HAS and BPE successfully extract more meaningful morphemes from more concatenative languages, despite all artificial languages being compositional. In the mixed concatenative and variable length languages, HAS yields lower metrics, suggesting it may indeed be less effective in languages with randomness. BPE can seemingly compensate for this as long as a sufficiently limited vocabulary size is chosen. The right vocabulary size is not obvious, however, as the BosDis for BPE is relatively low for artificial languages, suggesting an over-segmented language that erases the disentanglement of symbols. While segmentation algorithms behave as expected for systematic languages, these metrics are greater than 1 for random languages. This may result from over-sensitivity to noise in the absence of other signal, e.g., the presence of a few rare, highly predictive symbols may skew the metric upwards. This may also be attributed to a vanishingly small BoSDis for character-level segmentation. Future research should consider this edge case when comparing BoSDis before and after symbol segmentation.
@@
F-TopSim values indicate that the fusion language is indeed the most fusional. We observe that F-TopSim for random languages is zero, which does not suggest fusionality, rather that TopSim does not change with respect to fusion of attributes (expected for a random language, which should not assign any meaningful symbols to any attribute values or pairs thereof). We observe that fusionality does not vary significantly with respect to concatenativity, but increases in the variable length concatenative language.
@@
The evaluation metrics for these artificial languages are listed in Tables 4 and 5. Despite being compositional, these languages drastically impact evaluation metrics. The reordering and Mutation3 language are substantially less compositional (according to TopSim and BoSDis), segmentable, and learnable than fully concatenative languages. In Mutation0, where all characters are mutated twice, the languages were essentially indistinguishable from random languages by almost all metrics, including learnability. Even though this language consists entirely of messages formed by composing symbols according to simple, defined rules, in practice this compositionality may not matter or be recognized through existing metrics. The only metrics where there was a noticeable difference are the BoSDis ratios for HAS and BPE, but the interpretations of these are unclear, since segments cannot be disentangled symbols for this type of language.
@@
Appendix D HASLen for Natural and Emergent Languages

### v2

@@ 5.2 Natural Languages
As a comparison to humanlike communication schemes, we also analyze subsets of the verb conjugations of two natural languages, Spanish and Arabic, as Attr-Val languages. We chose these due to the availability of high-quality datasets, the fact that both conjugate for tense and person, and the fact that they represent fairly divergent communication schemes along the concatenativity axis. Spanish data comes from Jehle (2011), and Arabic data from Nawaz et al. (2025). We select a set of verb attributes comparable to the inflection Attr-Val setting: root, tense, and person. We then generate Attr-Val configurations based on combinations of these verb attributes sampled from the data, which serve as cognitively plausible (albeit simplified) representations of actual cognitive objects and events encoded and reconstructed in natural communication. For other attributes beyond tense and person, we choose a constant value (e.g., mood is always indicative). We generate 50 such inflectional sublanguages for Spanish and Arabic, each of which are mappings between lexeme-slot tuples and surface forms Wu et al. (2019).
Up to 76% of generated Spanish sublanguages can be meaningfully segmented by either HAS or BPE (i.e., BoSDis > 1), while only up to 8% of the Arabic sublanguages can be999Detailed analysis in Appendix G., demonstrating how both tools for symbol segmentation are insufficient for nonconcatenative morphology. Other metrics for these sublanguages are discussed in Section 6.
6 EmCom Experiments
Our Attr-Val settings targeting double articulation and inflection, general and specific evaluation metrics for communication schemes, and illustrative artificial and natural language reference points enable us to ask previously unexplored research questions around morphology in EmCom. To begin this exploration, we specifically ask:
@@ 6 EmCom Experiments
Do emergent languages exhibit fusion in inflection similarly to natural languages?
We next introduce experimental details, including the above articulation pressure. Guided by these questions, we then present the results.
6.1 Experimental Design
@@
Prior EmCom experiments have often explored pressures inspired by human evolution, such as ease of learning Kirby et al. (2014); Li and Bowling (2019) among others Vithanage et al. (2023); Galke and Raviv (2024), to encourage compositionality in emergent communication schemes. Meanwhile, prior work often assumes or favors concatenativity in evaluation Resnick et al. (2020); Ueda et al. (2023) despite the possibility of compositional but nonconcatenative communication schemes (e.g., our artificial and natural reference languages).
@@
In order to achieve concatenative emergent languages with double articulation, we explore an ease of articulation pressure. Human spoken languages are universally subject to phonological rules which determine the possibility of vocal sounds based on the environment and vocal tract physiology Dediu et al. (2017). The interface between phonology and morphology is a key area of study in linguistics, and a computational evolutionary model of morphology (as we hope to lay the groundwork for) would be incomplete without considering that morphemes are made up of meaningless components subject to similar rules independent of their meanings. Specifically, there are disallowed clusters of sounds in spoken languages (e.g., Hawaiian disallows consonant clusters), which may represent a significant pressure towards concatenative languages. To test this, we added a term to the training loss for a toy simulated phonology for agents, where even-valued characters can not occur next to even-valued characters, and odd-valued characters can not occur next to odd-valued characters.101010More details about this pressure provided in Appendix I.
In all experiments, agents are trained with 8 different random initializations. Learned languages are considered successful if the agent achieves 90% accuracy on the task, which occurs in most cases.1
| Setting | Art. | TopSim | BoSDis | HAS BoSDis | BPE96 BoSDis | BPELen96 | F-TopSim |
| ✗ | .339.034 | .213.047 | 3.254.31 | 1.25.359 | 3.82.235 | -.0418.0210 | |
| ✓ | .330.028 | .196.076 | 5.988.53 | 1.04.359 | 3.534.133 | -.0691.0350 | |
| ✗ | .247.119 | .504.0629 | 1.79.0943 | 1.23.0604 | 3.257 .225 | -.0078.028 | |
| ✓ | .144 0.102 | .509.0861 | 1.421.03 | 1.56 1.79 | 2.34.182 | -.0083.0178 | |
| ✗ | .0859 .039 | .6585 0.593 | 11.96.07 | 7.23 2.82 | 2.69 0.229 | .0415.0289 | |
| ✓ | 0.146.0860 | .3230.103 | 8.79 10.7 | 11.2 8.04 | 2.33 .226 | .0219.0185 | |
| ✗ | .0880.0367 | .582.136 | 2.17 6.07 | 1.472.82 | 2.690.212 | N/A | |
| ✓ | .239 .0645 | .645 .452 | 2.16 .775 | 2.51 1.25 | 2.330.181 | N/A |
6.2 Experimental Results
In Table 3, we summarize concatenativity and degree of fusion metrics in each Attr-Val setting, without and with articulation pressure. As shown there, BPE96 BoSDis > 1 on average across all emerged languages, implying the existence of multi-character meaningful segments and confirming that double articulation is achieved.
6.2.1 Concatenativity and Articulation
@@
We visualize the concatenativity (measured by BPELen) of emergent communication schemes in the default and inflection Attr-Val settings respectively in Figures 3 and 4. First, we observe no inductive bias toward concatenativity in our RNN-based agents as previous works implicitly assume; instead, emergent languages fall along a spectrum between perfectly concatenative and nonconcatenative. This may not be unrealistic: the natural languages also exhibit such a spectrum. Spanish appears relatively diverse, likely due to the presence of vowel patterns and auxilliary verbs. Arabic is more bimodal, with some very nonconcatenative sublanguages and some which have a high maximal compression due to being very fusional.
@@
However, the introduction of articulation pressure sharply decreases BPELen in emergent languages. We perform additional -tests on the mean BPELen96 of emergent languages between conditions, observing statistical significance in this change for both the default and inflection Attr-Val setting ( and respectively). This suggests ease of articulation is indeed a strong pressure towards concatenation, and the presence of “unpronounceable” clusters at the phonological (character) level is sufficient to induce concatenation at the morphological (symbol) level. More broadly, this may help explain the dominance of concatenative morphology in natural languages.
6.2.2 TopSim in Inflectional Languages
@@
To shed more light on the compositionality of naturalistic communication schemes, we compare the TopSim of emergent and natural languages in the inflection Attr-Val setting in Figure 5, including emergent languages with the articulation pressure. Compared to natural languages, TopSim for emergent languages was low, except for one instance with a TopSim of .507 (in the range of natural languages). In the 42 2 2 2 and 42 6 settings, but not in the 42 2 3 setting, the articulation pressure resulted in a significant increase in TopSim. There seems to be a complex interplay between this constraint and the specific attribute-value setup, which we leave for future work to explore in more detail. Together, this suggests that emergent languages still have room for improvement in achieving humanlike topographic similarity, and there may remain more pressures toward this to identify and explore.
6.2.3 Fusion Under Inflection
. indicating an inconsistent degree of fusion.
As we would predict, the 16x16x16 setting was least fusional, with a quite negative F-TopSim . In the 42 2 3 setting, F-TopSim was near 0, with 0 within a standard deviation of the mean. This implies there was some degree of fusion within these languages, since a completely non-fusional language would suffer a large drop in TopSim by fusing two attributes. The 42 2 2 2 setting was quite fusional, with a mean F-TopSim of 0.0415, i.e., by fusing the right attributes, we get a TopSim increase of over 0.04.)
@@ 6.2.3 Fusion Under Inflection
Suppletion (fusion of a root with grammatical features, e.g., went as the past tense of go) is much rarer than fusion between grammatical features. Our EmCom results mirror this: the pair of lowest-cardinality attributes (representing tense and person) had the highest F-TopSim in 6 of 8 of the 42 2 3 emergent languages, suggesting these attributes are most likely to fuse. Similar results were observed for the 42 2 2 2 setting. This may be explained by symbolic complexity Zhang (2025) increasing the least in fusing these attribute pairs; while 5 symbols are required to communicate the values for these two attributes before fusion, only 6 are required after. Meanwhile, fusing the root attribute (with 42 values) with another attribute can drastically increase the symbolic complexity (i.e., 2-3 times). This trend mirrors the fusion of fixed grammatical features in natural language, which we also observed in most of our generated Arabic and Spanish sublanguages.
@@ 7 Conclusion
In this work, we outline an experimental paradigm for emergent morphology, motivated by the principle of double articulation and a rich analogy between the Attr-Val game and inflectional morphology. We review existing metrics for EmCom, proposing new ones for concatenativity and degree of fusion. We calibrate our metrics towards a wide array of idealized compositional schemes that are possible under this paradigm, as well as schemes based on natural language inflection. We then demonstrate through EmCom experiments the power of this reimagined setting to produce linguistically interesting analyses grounded in comparison to natural language. In particular, the double articulation aspect of our paradigm reveals simulated articulatory constraints as a promising pressure toward concatenativity, which prior work has often implicitly assumed in evaluations of compositionality and attempted to elicit through other pressures, and future work may use to test more specific morphophonological hypotheses. Further, our emergent languages rederive the bias towards fusion of grammatical features in natural language. Future work should adopt small-vocabulary and inflectional Attr-Val settings for focused, linguistically motivated studies of morphology allowing direct comparison to natural language.
@@
Naturalness of task inputs.
Like related works Resnick et al. (2020); Chaabouni et al. (2020); Ueda et al. (2023), we chose to focus on Attr-Val reconstruction as our EmCom task setting. This setting removes the requirement of agents to perceive key features from stimuli (e.g., images), instead only requiring agents to learn to communicate ground truth features, significantly simplifying the learning problem. However, we intentionally chose this setting to ensure comparability with Ueda et al. (2023), and because our work is entirely focused on how EmCom agents learn to communicate features, thus possible perception errors would needlessly obstruct this effort. Further, considering that we aim to explore inflectional morphology in EmCom, the Attr-Val game is a natural analogy to existing models of morphology in linguistics, e.g., the lexeme-slot encoding of inflectional morphology Wu et al. (2019). Nonetheless, while out of scope for this work, an interesting direction for future work would be to investigate relationships between aspects of visual stimuli and morphology in emergent languages with small vocabularies.
@@
In our natural language reference points, we only considered Arabic and Spanish, which are not a representative sample of the world’s natural languages, and were chosen in part due to the availability of data. Without a much larger and more diverse dataset, we cannot definitively say whether a given emergent language is naturalistic. Additionally, the Attr-Val scheme we focused on is a rather small subset of the complexity of natural morphologies. A full analysis of a large set of natural languages, to cover this diversity of morphologies, is outside the scope of this paper, but we believe our examples will provide a solid initial point of comparison for emergent languages. Boldt and Mortensen (2024a) explored larger-scale cross comparisons of emergent languages, and we anticipate that the addition of parallel natural language data to such analyses would prove useful for future work.
@@
Appendix A Agent Implementation Details and Task Accuracy
In line with prior related works, we implement the sender and receiver agents as single layer gated recurrent unit (GRU; Cho et al., 2014) models using the EGG framework Kharitonov et al. (2019), with hidden dimensions of 500. Observed states were implemented by concatenating one-hot vectors for each attribute. Senders are optimized using REINFORCE Williams (1992) and receivers with backpropagation following Chaabouni et al. (2019) using the accuracy (i.e., proportion of attributes correctly reconstructed) as a loss, with cross-entropy loss as a supplementary loss term. It was found in initial experiments that this allowed smoother and more reliable convergence with lower sequence lengths, which is desirable for obtaining interpretable and minimally complex languages. Each run occurred on a single A40 GPU with 48GB VRAM and took about 8 hours, although many models converged before then. Models generally converged after about 400 training epochs.
We ran experiments with agents 8 times for each setting with different random initializations. In the default Attr-Val setting, 7 of 8 emerged languages were successful, i.e., achieved greater than 90% accuracy on the reconstruction task. For other experiments, including with the articulation pressure discussed in Section 6.2.1 and the inflectional Attr-Val setting, all languages were considered successful.
Appendix B Additional Notes On Inflection Setting
Appendix C Harris’ Articulation Scheme
We replicated the results in Ueda et al. (2023) that C-TopSim is greater than W-Topsim. This result led the authors to suggest that Criterion 3 for Harris’ Articulation Scheme (i.e., W-TopSim should be higher than C-TopSim) may be unsolvable.
@@ Appendix C Harris’ Articulation Scheme
Appendix D Positional Disentanglement
@@
Appendix E Artificial Languages Extended Analysis
@@
First, the level of concatenation in the first three artificial languages has minimal effect on TopSim and learnability, implying a lack of inductive bias toward concatenation in recurrent listener neural networks. Notably, this is not consistent with human results in linguistics Finley and Newport (2021), and implies additional pressures may be required to induce concatenative morphology. Meanwhile, as expected, BoSDis metrics indicate that both HAS and BPE successfully extract more meaningful morphemes from more concatenative languages, despite all artificial languages being compositional. In the mixed concatenative and variable length languages, HAS yields lower metrics, suggesting it may indeed be less effective in languages with randomness. BPE can seemingly compensate for this as long as a sufficiently limited vocabulary size is chosen. The right vocabulary size is not obvious, however, as the BosDis for BPE is relatively low for artificial languages, suggesting an over-segmented language that erases the disentanglement of symbols. While segmentation algorithms behave as expected for systematic languages, these metrics are greater than 1 for random languages. This may result from over-sensitivity to noise in the absence of other signal, e.g., the presence of a few rare, highly predictive symbols may skew the metric upwards. This may also be attributed to a vanishingly small BoSDis for character-level segmentation. Future research should consider this edge case when comparing BoSDis before and after symbol segmentation.
@@
F-TopSim values indicate that the fusion language is indeed the most fusional. We observe that F-TopSim for random languages is zero, which does not suggest fusion, rather that TopSim does not change with respect to fusion of attributes (expected for a random language, which should not assign any meaningful symbols to any attribute values or pairs thereof). We observe that degree of fusion does not vary significantly with respect to concatenativity, but increases in the variable length concatenative language.
@@
The evaluation metrics for these artificial languages are listed in Tables 5 and 6. Despite being compositional, these languages drastically impact evaluation metrics. The reordering and Mutation3 language are substantially less compositional (according to TopSim and BoSDis), segmentable, and learnable than fully concatenative languages. In Mutation0, where all characters are mutated twice, the languages were essentially indistinguishable from random languages by almost all metrics, including learnability. Even though this language consists entirely of messages formed by composing symbols according to simple, defined rules, in practice this compositionality may not matter or be recognized through existing metrics. The only metrics where there was a noticeable difference are the BoSDis ratios for HAS and BPE, but the interpretations of these are unclear, since segments cannot be disentangled symbols for this type of language.
@@
Appendix F HASLen for Natural and Emergent Languages